OpenStudy (anonymous):

Which of the following is a factor of 5y^4+4y^3+10y+8? a. y-5 b.5y-4 c. y^3+2 d. y^2-3 Thanks!!

OpenStudy (whpalmer4):

Do you know about the Rational Root Theorem?

OpenStudy (anonymous):

No sir/ma'am I do not.

OpenStudy (whpalmer4):

Well, when you multiply polynomials, the trailing component (\(8\), in this case) and the leading component (\(5y^4\), in this case) are "built up" exclusively from the trailing and leading components of everything you multiplied together. \[(x+2)(x+3) = x^2+2x+3x+2*3=x^2+5x+6\]6 comes from multiplying 2 and 3 \[(x+1)(x-4)=x^2-4x+1x+(-4)(1) = x^2-3x-4\]-4 comes from multiplying 1 and -4 \[(x+1)(x-4)(x+2) = (x^2-3x-4)(x+2) = \]\[= x^3+2x^2-3x^2-6x-4x-8 = x^3-x^2-10x-8\]-8 comes from multiplying 1, -4, and 2

OpenStudy (whpalmer4):

what this means is that if we are trying to figure out the factors of \(5y^4+4y^3+10y+8\), we can make some intelligent guesses about possible factors by factoring 8 and \(5y^4\).

OpenStudy (whpalmer4):

What are the factors of 8?

OpenStudy (anonymous):

The factors of 8 are 1, 2, 4, and 8

OpenStudy (whpalmer4):

right, so that means our choices have to include one of those numbers (or the negative of one of those numbers). For example, there's nothing we can multiply by \((y-5)\) to end up with a polynomial that ends with \(+8\).

OpenStudy (anonymous):

So I then start multiplying each option by the factors of 8?

OpenStudy (anonymous):

I think i figured out the answer. The answer is 5y-4.

OpenStudy (whpalmer4):

Mmm...don't think so...what do you multiply with \( (5y-4)\) to get \(5y^4+4y^3+10y+8\) ?

OpenStudy (whpalmer4):

The RRT lets us rule out some possible answer choices (a and d) but we still have a few to choose from (b and c).

OpenStudy (anonymous):

It would be y^3+2

OpenStudy (whpalmer4):

yes. \[5y(y^3+2) = 5y^4 + 10y\]so we need an additional \[5y^4+4y^3+10y+8-(5y^4+10y+8) = 4y^3+8\]and we can get that by \(4(y^3+2)\) so our factoring would be\[(5y+4)(y^3+2) = 5y^4+4y^3+10y+8\]

OpenStudy (whpalmer4):

If we had tried \(5y-4\) as a factor, we could get a \(y^3\): \[y^3(5y-4) = 5y^4-4y^3\]leaving us with \[5y^4+4y^3+10y+8-(5y^4-4y^3) = 8y^3+10y+8\]and unfortunately we can't factor out \(5y-4\) from that, so that would be a dead end. Make sense?

OpenStudy (whpalmer4):

Essentially, I used the RRT to eliminate candidates that couldn't possibly work, and then trial division to try the others.

OpenStudy (anonymous):

Yes! That makes much more sense now. Thank you very much!! :)

OpenStudy (whpalmer4):

There's another technique called grouping which would also work here: \[5y^4+4y^3+10y+8 = ( 5y^4 + 10y) + (4y^3 + 8)\]Now we factor each of the groups: \[5y(y^3+2) + 4(y^3+2)\]Now notice that we've factored the same thing out (but still in parentheses, instead of being out in front) in both terms, so we can rewrite that as \[(5y+4)(y^3+2)\]

OpenStudy (whpalmer4):

Then it's just a matter of looking for the matching answer.